OpenStudy (anonymous):

Well boys and girls its time for the first Chad's Tutorial of 2013 Sooooooo Here we goooooooooooooooooooooooooo.....

OpenStudy (anonymous):

Today we will learn how to add subtract divide and search google.

OpenStudy (anonymous):

Oh Yay!

OpenStudy (anonymous):

add: 2+2=4

OpenStudy (anonymous):

subtract: 2-2=0

OpenStudy (anonymous):

divide: 2 / 2 = 1

OpenStudy (anonymous):

Search google: type in box then press search

OpenStudy (anonymous):

Hope this helps :)

OpenStudy (anonymous):

Next time ill teach you guys how to multiply :)

OpenStudy (anonymous):

Oh Yesh!!!

OpenStudy (anonymous):

@KingGeorge @AccessDenied @karatechopper @saifoo.khan

OpenStudy (karatechopper):

Thats not..HOW to add, you are basically just showing the signs used in those operations.

OpenStudy (anonymous):

Just be able I tagged you I dont want your input

OpenStudy (anonymous):

Happy*

OpenStudy (karatechopper):

wha..

OpenStudy (anonymous):

Lol